The following sample of weights was taken from 9 cans of soda off the assembly line. Construct the 80% confidence interval for the population standard deviation for all cans of soda that come off the assembly line. Round your answers to two decimal places. 0.8,1.8,1.8,1.9,1.1,1.6,1.8,1.8,1.4
Answer #1

We have for given data,                                              
                                              
Sample variance =0.1453                                      
Sample standard deviation=0.3812                              
Sample size =9                                      
                                              
Degree of freedom =n-1=8                                  
Level of significance given =1-0.80=0.2                                  
Chi square critical value for lower tail=3.49                              
                                              
Chi square critical value for upper tail=13.362                              
                                              
                                          
We can use here confidence interval formula for population standard deviation                                              
(n-1) S2 (n-1).S2 V X1-a/2) V Xa/2)

(9 - 1) * 0.1453 13.362 (9 - 1) * 0.1453 3.49

=(0.295,0.577)   
                  
                                              
Lower confidence limit for population standard deviation=0.30   
                                              
Upper confidence limit for Population standard deviation=0.58